OpenStudy (anonymous):

What is the solution to the rational equation: 3/2 - 1/x-5 = -2/2x-10 x = -5 x = -5 over 2 No solution x = 5

OpenStudy (anonymous):

\[\frac{ a }{ b } +\frac{ c }{ d } =\frac{ ad+bc }{ bd }\] and also u have to know this \[\frac{ x }{ y }=\frac{ z }{ w } --->x \times w=y \times z\]